# Softpick: No Attention Sink, No Massive Activations with Rectified Softmax

### v1

@@ Abstract
We introduce softpick, a rectified, not sum-to-one, drop-in replacement for softmax in transformer attention mechanisms that eliminates attention sink and massive activations. Our experiments with 340M parameter models demonstrate that softpick maintains performance parity with softmax on standard benchmarks while achieving 0% sink rate. The softpick transformer produces hidden states with significantly lower kurtosis (340 vs 33,510) and creates sparse attention maps (46.97% sparsity). Models using softpick consistently outperform softmax when quantized, with particularly pronounced advantages at lower bit precisions. Our analysis and discussion shows how softpick has the potential to open new possibilities for quantization, low-precision training, sparsity optimization, pruning, and interpretability. Our code is available at https://github.com/zaydzuhri/softpick-attention.
Figure 1: (left) Comparison between the attention maps when using softmax vs softpick and overall sink rate of the models. (right) Largest hidden state activation per layer of the models
@@ 1 Introduction
The softmax function is widely used in statistics and particularly in machine learning as a way to normalize a vector of real numbers into a probability distribution. It has since been adopted as the de facto function to normalize the scores in the attention mechanism [Bahdanau et al., 2014, Sukhbaatar et al., 2015] used in the transformer architecture [Vaswani et al., 2017]. The use of softmax in attention was a natural choice, since it represents the probability of a query "matching" to a key among many keys, which can be used to return values each weighted by that probability. However, our modern use of attention makes us question this intuition [Miller, 2023, Smith, 2025]: why must they be non-zero probabilities that sum to one?
We continue to use softmax because in practice it is still the most capable normalization for attention. Softmax has great training stability with a neat Jacobian matrix resulting in a dense gradient, bounded Frobenius norm, and built-in regularization [Saratchandran et al., 2024], and has been proven to be sufficiently expressive in its nonlinearity [Chiang, 2024]. Despite the benefits, softmax brings several oddities detrimental to language modeling. Its sum-to-one nature forces out a now well-observed behavior named attention sink [Xiao et al., 2023] where attention heads allocate a significant score towards a specific token, often the initial BOS token, which itself is semantically irrelevant [Gu et al., 2024]. These sinks are, as far as we know, harmless to downstream performance. However, another symptom of this same effect of softmax is the appearance of extreme hidden state activations, dubbed massive activations [Sun et al., 2024], which grow larger as the model scales [Bondarenko et al., 2023]. These activations are especially problematic for quantization, with entire algorithms built around their existence. Additionally, most of the current low-precision training approaches try to work around these massive activations as they are a hassle to represent in low-bit precision. Removing these massive activations would open up a lot more possibilities.
@@
    We experiment by training 340M parameter transformer models from scratch and show that softpick achieves the same performance as softmax in benchmarks and training behavior. We also show that quantized softpick models down to 2-bit precision consistently outperform softmax.
@@ 2 Background
The attention sink as a phenomenon was first given a name in [Xiao et al., 2023] after observing that strong attention scores were given to tokens that were semantically irrelevant. These tokens can be any token that has minimal influence and appears often and consistently in training. Usually, the BOS token at the start of a sequence satisfies these conditions the best. Many works since then have observed, analyzed, and tried to explain when and why they appear. Barbero et al. [2025] argue that attention sinks are a way for LLMs to avoid over-mixing, where the attention heads that have a sink act as approximate no-ops as to avoid representational collapse in the residual stream of deep models. Meanwhile, in a more empirical work, Gu et al. [2024] investigate whether optimization, data distribution, loss function, or model architecture in language model pre-training is more correlated to the phenomenon. They conclude that attention sink happens for any transformer language model trained effectively on sufficient data. While optimizer hyperparameter choice can encourage a stronger sink, and loss function or data distribution can determine where the sink happens, ultimately the existence of the sink boils down to specifically the dependence of tokens’ attention scores to sum to one in the softmax normalization.
@@
Massive activations are so called in [Sun et al., 2024] to describe extremely large values in the hidden states or residuals of the transformer. Very few token positions actually reach these large magnitudes in some feature dimensions of their hidden states, where normally they are relatively much smaller in magnitude. These massive activations are important to consider because they hinder quantization and low-precision training to some degree [Dettmers et al., 2022]. Importantly, massive activations have been determined to be caused directly by self-attention and the softmax function inside it that manifests itself in the form of attention sink [Sun et al., 2024]. Therefore, getting rid of attention sink also gets rid of massive activations.
@@
Several of the studies mentioned above have attempted to mitigate attention sink to varying success. Gu et al. [2024] used sigmoid attention without normalization to avoid attention sink. However, attention sinks still appear if normalization is reintroduced, leading them to the conclusion that softmax is indeed the issue. They also experimented with adding attention bias by appending an extra vector to the queries, keys, and/or values in the sequence dimension. This was also tested in [Sun et al., 2024]. However, this only works if the vectors are trainable, which means adding extra parameters to the model. They also seem to worsen the deeper the model, as the massive activations appear again in later layers. Most promising is the approach in [Kaul et al., 2024] which tests an idea previously proposed in a blogpost [Miller, 2023], named softmax-1, adding a +1+1 in the denominator to relax the sum-to-one nature of softmax. However, this seems to only work well when used in conjuction with a custom optimizer, OrthoAdam, in order to avoid outlier activations. We propose softpick as a drop-in solution to mitigate both attention sink and massive activations at once, without the need for extra parameters or custom optimizers.
@@ 3.2 Design Rationale
The reason why we still do not modify this function too much is that the characteristics of softmax in training are ideal for attention. We want to maintain as much of the Jacobian matrix as possible and keep the gradient norm bounded that way. The next step would be to induce "null attention", or allowing attention to output zeros. We can achieve this by moving down the exponential by one, making it output negative values 1<y<0-1<y<0 for negative inputs. We then rectify this output with ReLU.
@@ 3.3 Derivative
with $step(x)$ being the step function that returns 0 if $x\leq 0$ and $1$ if $x>0$ and $sign(x)$ being the sign function that returns 1-1 if $x<0$ and $1$ if $x\geq 0$. Additionally, $\Sigma=\sum_{k=1}^{N}|e^{x_{k}-m}-e^{-m}|+\epsilon$ is the denominator of the softpick function. Although unlike softmax where only the outputs need to be kept for a naive backward pass implementation, the softpick derivative is nonetheless trivial for implementations that recompute the inputs, as is done in FlashAttention.
@@ 3.4 FlashAttention
The FlashAttention algorithm [Dao et al., 2022, Dao, 2023] allows for online single-pass computation of attention, bypassing the quadratic memory requirement. Because the $ReLU(x)$ and absolute $|x|$ functions uphold the multiplicative property for positive-valued multipliers (such as $e^{x}$ for any $x$), deriving an online version of softpick is possible, thus making it compatible with FlashAttention. We provide the algorithms for the more recent FlashAttention-2 version of the forward pass and backward pass that use the softpick function instead of softmax in Appendix A.
@@ 4 Experiments
To validate the viability of the softpick function over softmax, we conduct experiments on two transformer models trained from scratch. Both are Llama-style (pre-norm, RoPE, SwiGLU MLP) transformers with 340M parameters (373M with embeddings), one using the vanilla softmax function in attention, and the other using the softpick function instead of softmax. The detailed training configuration for both models is shown in Table 1. We use the flash-linear-attention repository [Yang and Zhang, 2024] and their Flame training framework (built on top of torchtitan [Liang et al., 2025]) because of their easily modifiable but fast triton [Tillet et al., 2019] kernels. We train each model on a total of 52 billion tokens sampled from the 100B subset of the fineweb-edu [Lozhkov et al., 2024] dataset. Each training run took approximately 18 hours on 8xH100 GPUs.
Table 1: Training configuration and hyperparameters.
@@
Hidden size 1024 Optimizer AdamW
Num. layers 24 Learning rate 3e-4
Num. heads 16 LR schedule cosine
Num. KV heads 16 Minimum LR (%) 10%
Sequence length 4096 Warmup steps 1000
RoPE $\theta$ 10,000 Device batch size 16
@@
Vocab size 32,000 Training steps 100,000
@@ 5 Results
Figure 2: Training loss and gradient norm during training.
We present the resulting training loss and gradient norm over 100,000 training steps (52B tokens) in Figure 2. We observed that the training loss of the softpick transformer tracks very closely to the vanilla softmax transformer, with only a 0.004 gap in the final training loss. This supports our design decision to maintain as much of the important components of the softmax function as possible. We observe that the total gradient norm of the models’ parameters during training exhibits noteworthy behavior. At the start up to a third of the way into training, the gradient norm of the softpick model is higher than softmax, but goes down and stays below softmax until the end, and does not rise at as fast of a rate as softmax. Also important to note is that the magnitude of some gradient norms of the softpick model are much higher than the gradient norm peaks of softmax. We use gradient clipping with a maximum norm of 1.0 and do not see any instability caused by these large gradients. Overall, we maintain the favorable training properties of softmax with softpick.
@@
Table 2: Comparison of softpick vs softmax performance on downstream tasks. $\uparrow$ = higher is better, $\downarrow$ = lower is better. $\Delta$ = softpick - softmax.
Task Metric Softmax Softpick $\Delta$ (S - V)
Arc Easy Acc Norm $\uparrow$ 56.31 56.61 +0.30
Acc $\uparrow$ 60.61 60.82 +0.21
Lambada Acc $\uparrow$ 36.35 36.21 -0.14
Perplexity $\downarrow$ 30.33 28.67 -1.66
Piqa Acc Norm $\uparrow$ 66.43 66.27 -0.16
Acc $\uparrow$ 66.87 66.43 -0.44
Sciq Acc Norm $\uparrow$ 75.10 77.20 +2.10
Acc $\uparrow$ 83.30 83.60 +0.30
Wikitext Word Perplexity $\downarrow$ 204.94 141.56 -63.38
We benchmark both models with general downstream tasks that are appropriate for the model size. We use the LM Evaluation Harness [Gao et al., 2023] to evaluate on these benchmarks: AI2 Reasoning Challenge (ARC-e) [Clark et al., 2018], Lambada [Paperno et al., 2016], specifically the OpenAI variant, PIQA [Bisk et al., 2019], and SciQ [Welbl et al., 2017]. All 4 benchmarks report accuracy, with an addition of perplexity for Lambada. We also measure validation perplexity on Wikitext [Merity et al., 2016]. We observe equal if not slightly better performance from softpick compared to softmax. The perplexity on both Wikitext and Lambada is better (lower) than softmax, which is opposite to the training loss that is marginally higher for softpick. All in all, softpick does not detract from and in some cases might add to the performance of softmax in attention.
@@
Figure 3: Quantization result of softmax vs softpick.
We also benchmark the effect of quantization on the trained models. We verified this using HQQ [Badri and Shaji, 2023], BitsandBytes (BNB) [Dettmers et al., 2022, 2023], and GPTQ [Frantar et al., 2023]. These methods encompass different approaches, including those that utilize calibration data and those that do not. We use the same benchmarks as before. The results can be seen in Figure  3 and more complete tables in Appendix C. When using softpick, the quantized models are consistently better than when softmax is used, and quantization is less damaging as the precision goes down. We discuss why this is in more detail in 7.1 after analyzing massive activations in 6.3.
@@ 6 Analysis
Table 3: Metrics for analysis of softmax and softpick in terms of sink rate, hidden activation distribution, and attention map sparsity.
Model Sink Rate % Hidden Activation Sparsity %
$\epsilon_{s}=0.6$ $\epsilon_{s}=0.3$ Kurtosis Min. Max.
Softmax 33.99 63.41 33510.81 -207.55 240.27 0.00
Softpick 00.00 00.00 340.96 -45.03 36.21 46.97
Figure 4: Box plots of the hidden state activations at every 2 layers + last layer of the softmax (left, blue) and softpick (right, orange) models.
@@ 6.1 Attention Maps
Figure 5: Attention maps of softmax and softpick models on 3 different input texts. Two heads are visualized: Head 4 of Layer 10 and Head 9 of Layer 21. See more attention maps in Appendix D.
To understand the difference that the softpick function makes compared to softmax, it is best to observe the attention maps directly. Attention maps are the matrices resulting from the computation of $\mathbf{A}=Soft\{max,pick\}(\mathbf{QK}^{T}/\sqrt{d_{k}})$. We prepare three text inputs as samples to run through the models and extract the attention maps. We provide some clear example attention maps in Figure 5 and show more attention maps of the models at every 4 layers and every 4 heads on the 3 inputs in Appendix D.
@@
To see the sparsity effect of softpick, we calculate the sparsity ratio of the attention map for softpick and softmax by running 10 text samples through the model, collecting the attention maps, and averaging the results. Note that since this is a causal language model, we only consider the lower triangular scores of the attention map. As shown in Table 3, our method results in attention scores with 46.97% sparsity.

### v2

@@ Abstract
We introduce softpick, a rectified, not sum-to-one, drop-in replacement for softmax in transformer attention mechanisms that eliminates attention sink and massive activations. Our experiments with 340M and 1.8B parameter models demonstrate that softpick achieves 0% sink rate consistently. The softpick transformers produce hidden states with significantly lower kurtosis and creates sparse attention maps. Quantized models using softpick outperform softmax on standard benchmarks, with a particularly pronounced advantage at lower bit precisions. Our analysis and discussion shows how softpick has the potential to open new possibilities for quantization, low-precision training, sparsity optimization, pruning, and interpretability. Our code is available at https://github.com/zaydzuhri/softpick-attention.
Figure 1: (left) Comparison between the attention maps when using softmax vs softpick and overall sink rate of the 340M models. (right) Largest hidden state activation per layer of the 340M models.
@@ 1 Introduction
The softmax function is widely used in statistics and particularly in machine learning as a way to normalize a vector of real numbers into a probability distribution. It has since been adopted as the de facto function to normalize the scores in the attention mechanism [3, 43] used in the transformer architecture [46]. The use of softmax in attention was a natural choice, since it represents the probability of a query "matching" to a key among many keys, which can be used to return values each weighted by that probability. However, our modern use of attention makes us question this intuition [29, 41]: why must they be non-zero probabilities that sum to one?
We continue to use softmax in practice because it is still the most capable normalization for attention. Softmax has great training stability with a neat Jacobian matrix resulting in a dense gradient, bounded Frobenius norm, and built-in regularization [37], and has been proven to be sufficiently expressive in its nonlinearity [7]. Despite the benefits, softmax brings several oddities detrimental to language modeling. Its sum-to-one nature forces out a now well-observed behavior named attention sink [51] where attention heads allocate a significant score towards a specific token, often the initial BOS token, which itself is semantically irrelevant [19]. These sinks are, as far as we know, harmless to downstream performance. However, another symptom of this same effect of softmax is the appearance of extreme hidden state activations, dubbed massive activations [44], which grow larger as the model scales [6]. These activations are especially problematic for quantization, with entire algorithms built around their existence. Additionally, most of the current low-precision training approaches try to work around these massive activations as they are a hassle to represent in low-bit precision. Removing these massive activations would open up a lot more possibilities.
@@
    We experiment by training 340M and 1.8B parameter transformer models from scratch and show how softpick compares to softmax in benchmarks and training behavior. We also show that quantized softpick models down to 2-bit precision outperform softmax.
@@ 2 Background
The attention sink as a phenomenon was first given a name in [51] after observing that strong attention scores were given to tokens that were semantically irrelevant. These tokens can be any token that has minimal influence and appears often and consistently in training. Usually, the BOS token at the start of a sequence satisfies these conditions the best. Many works since then have observed, analyzed, and tried to explain when and why they appear. Barbero et al. [4] argue that attention sinks are a way for LLMs to avoid over-mixing, where the attention heads that have a sink act as approximate no-ops as to avoid representational collapse in the residual stream of deep models. Meanwhile, in a more empirical work, Gu et al. [19] investigate whether optimization, data distribution, loss function, or model architecture in language model pre-training is more correlated to the phenomenon. They conclude that attention sink happens for any transformer language model trained effectively on sufficient data. While optimizer hyperparameter choice can encourage a stronger sink, and loss function or data distribution can determine where the sink happens, ultimately the existence of the sink boils down to specifically the dependence of tokens’ attention scores to sum to one in the softmax normalization.
@@
Massive activations are so called in [44] to describe extremely large values in the hidden states or residuals of the transformer. Very few token positions actually reach these large magnitudes in some feature dimensions of their hidden states, where normally they are relatively much smaller in magnitude. These massive activations are important to consider because they hinder quantization and low-precision training to some degree [13]. Importantly, massive activations have been determined to be caused directly by self-attention and the softmax function inside it that manifests itself in the form of attention sink [44]. Therefore, getting rid of attention sink also gets rid of massive activations.
@@
Several of the studies mentioned above have attempted to mitigate attention sink to varying success. Gu et al. [19] used sigmoid attention without normalization to avoid attention sink. However, attention sinks still appear if normalization is reintroduced, leading them to the conclusion that softmax is indeed the issue. They also experimented with adding attention bias by appending an extra vector to the queries, keys, and/or values in the sequence dimension. This was also tested in [44]. However, this only works if the vectors are trainable, which means adding extra parameters to the model. They also seem to worsen the deeper the model, as the massive activations appear again in later layers. Most promising is the approach in [21] which tests an idea previously proposed in a blogpost [29], named softmax-1, adding a $+1$ in the denominator to relax the sum-to-one nature of softmax. However, this seems to only work well when used in conjuction with a custom optimizer, OrthoAdam, in order to avoid outlier activations. Owen et al. [33] further confirms that these KV bias techniques are not a reliable mitigation strategy for massive activations. We propose softpick as a drop-in solution to mitigate both attention sink and massive activations at once, without the need for extra parameters or custom optimizers.
@@ 3.2 Design Rationale
The reason why we still do not modify this function too much is that the characteristics of softmax in training are ideal for attention. We want to maintain as much of the Jacobian matrix as possible and keep the gradient norm bounded that way. The next step would be to induce "null attention", or allowing attention to output zeros. We can achieve this by moving down the exponential by one, making it output negative values $-1<y<0$ for negative inputs. We then rectify this output with ReLU.
@@ 3.3 Derivative
with $step(x)$ being the step function that returns 00 if $x\leq 0$ and $1$ if $x>0$ and $sign(x)$ being the sign function that returns $-1$ if $x<0$ and $1$ if $x\geq 0$. Additionally, $\Sigma=\sum_{k=1}^{N}|e^{x_{k}-m}-e^{-m}|+\epsilon$ is the denominator of the softpick function. Although unlike softmax where only the outputs need to be kept for a naive backward pass implementation, the softpick derivative is nonetheless trivial for implementations that recompute the inputs, as is done in FlashAttention.
@@ 3.4 FlashAttention
The FlashAttention algorithm [10, 9] allows for online single-pass computation of attention, bypassing the quadratic memory requirement. Because the $ReLU(x)$ and absolute $|x|$ functions uphold the multiplicative property for positive-valued multipliers (such as $e^{x}$ for any $x$), deriving an online version of softpick is possible, thus making it compatible with FlashAttention. We provide the algorithms for the more recent FlashAttention-2 version of the forward pass and backward pass that use the softpick function instead of softmax in Appendix A.
@@ 4 Experiments
To validate the viability of the softpick function over softmax, we conduct experiments on Llama-style (pre-norm, RoPE, SwiGLU MLP) transformers trained from scratch. We train 4 models, two for each size: 340M and 1.8B parameters, each with a softmax variant and a softpick variant. The detailed training configuration for all models is shown in Table 1. We use the flash-linear-attention repository [53] and their Flame training framework (built on top of torchtitan [23]) because of their easily modifiable but fast triton [45] kernels. We train each 340M model on a total of 52 billion tokens and each 1.8B model on 104 billion tokens sampled from the 100B subset of the fineweb-edu [25] dataset. Each training run took approximately 18 hours for the 340M models and 116 hours for the 1.8B models on 8xH100 GPUs.
Table 1: Training configuration and hyperparameters for {340M, 1.8B} models.
@@
Hidden size {1024, 2048} Optimizer AdamW
Num. layers {24, 32} Learning rate {3e-4, 2e-4}
Num. heads {16, 32} LR schedule cosine
Num. KV heads {16, 32} Minimum LR (%) 10%
Sequence length 4096 Warmup steps {1000, 2000}
RoPE $\theta$ 10,000 Device batch size {16, 8}
Grad. accum. steps {1, 2}
@@
Vocab size 32,000 Training steps {100000, 200000}
@@ 5 Results
Figure 2: Training loss and gradient norm during training of 340M models.
We present the resulting training loss and gradient norm of the 340M model over 100,000 training steps (52B tokens) in Figure 2. We observed that the training loss of the softpick transformer tracks very closely to the vanilla softmax transformer, with only a 0.004 gap in the final training loss. This supports our design decision to maintain as much of the important components of the softmax function as possible. However, this does not hold at scale, as we observe a much larger gap of 0.12 in the training loss of the larger 1.8B model. This will be heavily reflected in the benchmark results. Tangentially, we observe that the total gradient norm of the models’ parameters during training exhibits noteworthy behavior. At the start up to a third of the way into training, the gradient norm of the softpick model is higher than softmax, but goes down and stays below softmax until the end, and does not rise at as fast of a rate as softmax. Also important to note is that the magnitude of some gradient norms of the softpick model are much higher than the gradient norm peaks of softmax. We use gradient clipping with a maximum norm of 1.0 and do not see any instability caused by these large gradients.
@@
Table 2: Comparison of softpick vs softmax 340M parameter model performance on downstream tasks. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 56.61 56.73 +0.12
Acc $\uparrow$ 60.35 61.11 +0.76
Lambada Acc $\uparrow$ 36.25 36.21 -0.04
Perplexity $\downarrow$ 30.33 28.63 -1.70
Piqa Acc Norm $\uparrow$ 66.59 66.49 -0.10
Acc $\uparrow$ 66.97 66.59 -0.38
Sciq Acc Norm $\uparrow$ 74.90 77.30 +2.40
Acc $\uparrow$ 83.20 83.60 +0.40
Wikitext Word Perplexity $\downarrow$ 23.85 24.32 +0.47
Table 3: Comparison of softpick vs softmax 1.8B parameter model performance on downstream tasks. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 67.21 62.04 -5.17
Acc $\uparrow$ 72.73 68.60 -4.13
Lambada Acc $\uparrow$ 49.43 43.51 -5.92
Perplexity $\downarrow$ 11.38 15.81 +4.43
Piqa Acc Norm $\uparrow$ 73.61 70.89 -2.72
Acc $\uparrow$ 73.78 71.27 -2.51
Sciq Acc Norm $\uparrow$ 86.40 80.40 -6.00
Acc $\uparrow$ 90.10 87.00 -3.10
Wikitext Word Perplexity $\downarrow$ 15.10 17.87 +2.77
We benchmark all models with general downstream tasks that are appropriate for the model sizes. We use the LM Evaluation Harness [18] to evaluate on these benchmarks: AI2 Reasoning Challenge (ARC-e) [8], Lambada [34], specifically the OpenAI variant, PIQA [5], and SciQ [49]. All 4 benchmarks report accuracy, with an addition of perplexity for Lambada. We also measure validation perplexity on Wikitext [27]. See Table 2 and Table 3 for full results. We observe equal if not slightly better performance from softpick compared to softmax on the 340M model. However, the 1.8B models show that softpick does not seem to scale well, at least when using the same hyperparameters as softmax. Overall accuracy on benchmarks and perplexity are worse at this scale.
@@
Figure 3: Quantization result of softmax vs softpick 340M parameter models.
Figure 4: Quantization result of softmax vs softpick 1.8B parameter models.
We also benchmark the effect of quantization on the trained models. We verified this using HQQ [2], BitsandBytes (BNB) [13, 14], and GPTQ [17]. These methods encompass different approaches, including those that utilize calibration data and those that do not. We use the same benchmarks as before. The results can be seen in Figure 3 and Figure 4 with more complete tables in Appendix C. At the 340M scale, when using softpick, the quantized models are consistently better than when softmax is used, and quantization is less damaging as the precision goes down. This becomes less clear at the 1.8B as the gap in performance of the base models is larger. However, the gap does close in the lower the precision. We discuss why this is in more detail in 7.1 after analyzing massive activations in 6.3.
@@ 6 Analysis
Table 4: Metrics for analysis of softmax and softpick in terms of sink rate, hidden activation distribution, and attention map sparsity. *Note that sparsity was calculating by counting exact zeros in the lower triangle, which means the softmax sparsity only comes from underflowed values.
Size Model Sink Rate % Hidden Activation Sparsity %
$\epsilon_{s}=0.2$ $\epsilon_{s}=0.3$ Kurtosis Min. Max.
340M Softmax 68.28 63.41 33510.81 -207.55 240.27 4.53*
Softpick 00.00 00.00 340.96 -45.03 36.21 99.34
1.8B Softmax 41.73 14.96 74456.81 -173.48 371.43 4.28*
Softpick 00.00 00.00 2193.27 -51.05 101.51 95.63
(a) 340M models
(b) 1.8B models
Figure 5: Box plots of the hidden state activations at some layers of the softmax (left, blue) and softpick (right, orange) models.
@@ 6.1 Attention Maps
Figure 6: Attention maps of softmax and softpick 340M models on 3 different input texts. Two heads are visualized: Head 4 of Layer 10 and Head 9 of Layer 21. See more attention maps in Appendix E.
To understand the difference that the softpick function makes compared to softmax, it is best to observe the attention maps directly. Attention maps are the matrices resulting from the computation of $\mathbf{A}=Soft\{max,pick\}(\mathbf{QK}^{T}/\sqrt{d_{k}})$. We prepare three text inputs as samples to run through the models and extract the attention maps. We provide some clear example attention maps from the 340M models in Figure 6 and show more attention maps of all models in Appendix E.
@@
To see the sparsity effect of softpick, we calculate the sparsity ratio of the attention map for softpick and softmax by running 10 text samples through the model, collecting the attention maps, and averaging the the sparsity, which is calculated by counting the number of exact zeros. Note that since this is a causal language model, we only consider the lower triangular scores of the attention map. As shown in Table 4, our method results in attention scores with 99.34% sparsity on the 340M model and 95.63% on the 1.8B model. While softmax in theory cannot return exact zeros, we observe a non-zero sparsity because in practice floating point scores can underflow to zero.